OpenStudy (anonymous):

Solve 53x + 1 = 25x + 1

OpenStudy (anonymous):

5x+1=25x+1 20x=2 x=2/20 x=1/10

OpenStudy (anonymous):

@ywping How did you get 28x=2?

OpenStudy (anonymous):

The 1s cancels each other. x=0.

OpenStudy (anonymous):

@pastory It's 53x

OpenStudy (anonymous):

Sry for my error previously. 53x=25x+1-1 28x=0